### v3

# GeoAware-VLA: Implicit Geometry Aware Vision-Language-Action Model

Abstract
Vision-Language-Action (VLA) models often fail to generalize to novel camera viewpoints, a limitation stemming from their difficulty in inferring robust 3D geometry from 2D images. We introduce GeoAware-VLA, a simple yet effective approach that enhances viewpoint invariance by integrating strong geometric priors into the vision backbone. Instead of training a visual encoder or relying on explicit 3D data, we leverage a frozen, pretrained geometric vision model as a feature extractor. A trainable projection layer then adapts these geometrically-rich features for the policy decoder, relieving it of the burden of learning 3D consistency from scratch. Through extensive evaluations on LIBERO benchmark subsets, we show GeoAware-VLA achieves substantial improvements in zero-shot generalization to novel camera poses, boosting success rates by over 2x in simulation. Crucially, these benefits translate to the physical world; our model shows a significant performance gain on a real robot, especially when evaluated from unseen camera angles. Our approach proves effective across both continuous and discrete action spaces, highlighting that robust geometric grounding is a key component for creating more generalizable robotic agents. Project Page https://alisharey.github.io/GeoAware-VLA/
I Introduction
The development of general-purpose agents capable of performing diverse manipulation tasks in unstructured environments is a central goal in robotics. Vision-Language-Action (VLA) model is a popular paradigm, which learns to map visual observations and natural language instructions directly to robot actions. Most methods achieve strong performance within their training domains [haldar2024baku, kim2024openvla, zitkovich2023rt], but often fail to generalize beyond them, struggling even with minor changes in camera viewpoints [simpler, xie2024decomposing]. This limitation arises from the difficulty of inferring a consistent 3D world model from 2D visual inputs, a prerequisite for reliable manipulation and spatial awareness.
Prior work has explored two main strategies to address this challenge. One approach is to incorporate explicit 3D representations, such as point clouds, into the policy’s observation space [wilcox2025adapt3r, yang2025fp3, ze20243d]. While effective, these methods often require depth sensors and introduce significant computational overhead for constructing and processing explicit 3D structures. An alternative lies in implicitly encouraging the model to learn geometrically consistent features without reconstructing the scene in 3D. This is often achieved by leveraging multi-view data or data augmentation during training to produce a view-invariant policy [pang2025learning, tian2024view].
Figure 1: (Top) Illustration of the training and testing viewpoints in the LIBERO dataset. GeoAware-VLA demonstrates zero-shot generalization across novel viewpoints. (Mid) GeoAware-VLA achieves superior generalization to novel viewpoints, outperforming the state-of-the-art by over 30% in success rate on the LIBERO dataset. (Bottom) Key intermediate steps during successful real-world deployment.
However, these implicit methods also have their own drawbacks. Disentanglement techniques often rely on curated multi-view datasets for effective training, while augmentation-based methods are fundamentally limited by the computational cost of producing novel views and the distribution of views from which they are sampled. These challenges can be addressed by leveraging a powerful, pretrained geometric foundation models that have already distilled 3D understanding from vast datasets. For instance, the Visual Geometry Grounded Transformer (VGGT) [wang2025vggt] is explicitly optimized to distill a rich 3D representation of the world, having been trained to infer key geometric aspects such as camera parameters, multi-view depth, dense point clouds, and point tracking.
Built on these insights, we propose GeoAware-VLA, a simple yet highly effective modification to a standard VLA architecture that significantly enhances robustness to novel camera viewpoints. Our core hypothesis is that a policy’s ability to generalize across views is fundamentally tied to the geometric acuity of its visual encoder. We therefore replace the standard image encoder with the powerful VGGT backbone, using it as a frozen feature extractor. By building upon features that are already inherently consistent across different perspectives, our policy is relieved of the burden of learning 3D geometry from scratch. We need only append a lightweight, trainable projection layer to map these potent features into the latent space of the BAKU policy decoder [haldar2024baku]. As our experiments demonstrate, the integration of a powerful geometric prior results in substantial improvements in zero shot generalization to novel camera viewpoints as shown in Figure 1.
In summary, our contributions are as follows:
  • We propose GeoAware-VLA an effective integration of visual geometry foundational models into VLAs.
  • GeoAware-VLA achieves a 2x success rate increase on novel views and shows significant performance gains on a physical robot.
  • We confirm the versatility of our method by showing it consistently improves generalization across different policy decoders.
Figure 2: The overall architecture diagram of GeoAware-VLA. It inputs multi-view images, uses VGGT to extract view robust features to generate robot actions.
II Related Work
II-A Imitation Learning
Modern robotic learning is rooted in Imitation Learning (IL), where early methods like Behavioral Cloning (BC) [pomerleau1988alvinn] struggled with compounding errors from distribution shift. This motivated the development of more robust visuomotor policies, with generative methods like Diffusion Policy [chi2023diffusion] and energy-based models like Implicit Behavioral Cloning (IBC) [florence2022implicit] learning complex action distributions to enhance control. Building on this robustness, the field’s focus shifted towards creating large-scale, generalist agents. Influential models like DeepMind’s Gato [reedgeneralist] and Octo [ghosh2024octo] demonstrated that a single, large transformer pre-trained on diverse, multi-robot datasets could establish a powerful and adaptable motor control foundation. This trajectory culminated in the current paradigm of VLA models, which unify perception, language, and control by leveraging knowledge from web-scale models. This includes a series of influential works such as RT-1 [RT-1] and its successors [RT-2, RT-X], alongside powerful open-source models like OpenVLA [kim2024openvla, kim2025fine] and $\pi$ -[black2410pi0, intelligence2504pi05], which have collectively pushed the boundaries of generalization and performance. Our work builds on the BAKU architecture [haldar2024baku], a lightweight VLA that stands out for its efficient design choices, including multi-sensory observations and action chunking.
II-B Vision Encoders in Robotic Learning
The vision backbone, which processes raw pixel inputs, is a critical component of VLA models. Foundational approaches often used semantic encoders like ResNet [he2016deep], and later enhanced with FiLM layers to incorporate conditioning from language or other inputs [perez2018film].
A more recent paradigm involves models like SigLIP [zhai2023sigmoid] and its successor SigLIP 2 [tschannen2025siglip], which are pre-trained on vast web-scale image-text pairs. This training method creates representations that are naturally aligned with language, making them highly effective for VLA tasks. However, the primary focus of these backbones remains semantic - identifying what objects are, rather than their precise 3D geometry.
Alongside the development of semantic encoders for policy learning, a distinct class of powerful geometric vision models has emerged from the computer vision community. Architectures such as DUSt3R [wang2024dust3r], MUST3R [cabon2025must3r], and VGGT [wang2025vggt] are systems designed specifically for geometric inference from multi-view imagery, inferring properties like camera poses and dense depth maps. While highly effective for 3D reconstruction, these models were not intended as general-purpose backbones for policy learning and, consequently, their use in robotics has been limited. To our knowledge, such geometric models have not been implicitly integrated into VLA models before. Our work provides an initial exploration in this direction, investigating how to bridge this gap by leveraging the implicit 3D understanding from these geometric models to improve robotic policies.
II-C View-point robust robotic policies
To achieve more viewpoint-robust policies, various methods have been explored. One approach involves augmenting training data with images rendered from simulators [pang2025learning, seo2023multi]. While effective, this introduces the challenge of sim-to-real transfer. Another promising way is utilizing novel view synthesis [tian2024view, yang2025mobi]. Unlike simulator-based methods, these models can operate directly on real-world data, circumventing the sim-to-real gap. However, these models may still perform poorly when the new viewpoint is drastically different from the original view [tian2024view].
Other works, like Act3D [gervet2023act3d] and 3D Diffuser Actor [ke20243ddiffuser], utilize explicit 3D representations such as point clouds or voxel grids [shridhar2023perceiver, zhulearning, ze20243d, wilcox2025adapt3r]. These methods are robust to viewpoint changes but often require reliable depth data and camera calibration, which can be scarce in large-scale datasets like [RT-X, libero, walke2023bridgedata]. In contrast to these methods, our approach integrates strong geometric priors, which relieves the policy of the burden of learning 3D consistency from scratch and avoids the need for explicit depth or calibration information.
III Problem Setup
We consider a standard setup for multi-task imitation learning where the goal is to train a policy $\pi(a_{t}|o_{t},l)$ that predicts a robot action $a_{t}$ given the current observation $o_{t}$ and a natural language instruction $l$. The observation $o_{t}$ consists of a set of $V$ RGB images from different camera viewpoints, denoted as $\mathbf{I_{t}}=\{I_{1},I_{2},\dots,I_{V}\}$, and a robot proprioceptive state $\mathbf{p}_{t}$. Together, these form the complete observation $o_{t}=(\mathbf{I_{t}},\mathbf{p}_{t})$. The policy is parameterized by a neural network $\pi_{\theta}$ with trainable weights $\theta$. The action $a_{t}\in\mathbb{R}^{7}$ at each timestep $t$ specifies the relative end-effector pose and the gripper position. This 7-dimensional action vector is defined as:
$a_{t}=[\Delta x,\Delta y,\Delta z,\Delta\alpha,\Delta\beta,\Delta\gamma,g_{t}]$
where:
  • $(\Delta x,\Delta y,\Delta z)$ is the 3D vector for the relative translation of the end-effector.
  • $(\Delta\alpha,\Delta\beta,\Delta\gamma)$ is the 3D axis-angle representation for the relative rotation of the end-effector.
  • $g_{t}\in\{+1,-1\}$ is a scalar representing the gripper state, where +1+1 signifies ’open’ and 1-1 signifies ’closed’.
The policy is trained on a dataset of expert demonstrations $\mathcal{D}=\{(o_{t}^{(i)},l^{(i)},a_{t}^{(i)})\}_{i=1}^{N}$ using Behavior Cloning (BC), which frames the imitation learning task as a supervised learning problem. The objective is to find the optimal parameters $\theta^{*}$ for the policy network $\pi_{\theta}$ that produce actions as close as possible to the expert actions $a_{t}$ from the demonstration dataset $\mathcal{D}$. This is achieved by minimizing a loss function, such as the Mean Squared Error (MSE), with respect to the network parameters $\theta$:
[…]
Here, the expectation $\mathbb{E}$ is taken over all samples in the expert demonstration dataset $\mathcal{D}$. By minimizing this loss, the policy network learns to map observations and language commands to the corresponding expert-level actions.
IV Methodology
We propose GeoAware-VLA, a novel approach to enhance a standard VLA framework. The core of our method is the replacement of the trainable vision encoder with a frozen, Geometrically-Aware one. To make the rich, multi-layer features from the VGGT backbone usable by the policy network, we introduce a trainable vision projection layer. The remainder of the policy architecture adapts and builds upon elements from [haldar2024baku]. The overall architecture of our model is illustrated in Figure 2
The overall architecture can be divided into three stages: sensory encoding, policy decoding, and action generation. The final action $a_{t}$ is a function of the visual observation $o_{t}^{\text{vis}}$, proprioceptive state $s_{t}$, and language instruction $l_{t}$:
$a_{t}=\pi_{\phi}(P_{\theta}(E_{\text{VGGT}}(o_{t}^{\text{vis}})),E_{\text{lang}}(l_{t}),E_{\text{proprio}}(s_{t}))$
where $E$ represents the respective encoders, $P_{\theta}$ is our trainable vision projection layer, and $\pi_{\phi}$ is the policy network comprising the transformer trunk and action head.
IV-A Sensory Encoders
Each input modality is processed by a specialized encoder to generate a fixed-dimension embedding, $D_{\text{repr}}$, before being passed to the policy’s observation trunk.
Figure 3: A visual representation of a single episode from our experimental setup, which shows how different viewpoints (rows) change over time (left to right columns). The top row shows the wrist camera images, while the second row displays the top-down viewpoint. The model is trained exclusively on these two viewpoints. The bottom three rows show three novel, unseen viewpoints used for evaluating the model
IV-A1 Geometrically-Aware Vision Encoder
To encode visual observations, we utilize a frozen, pretrained VGGT model backbone and remove the prediction heads. A key property of VGGT is that it produces a list of feature tensors from multiple intermediate layers, capturing a hierarchy of visual and geometric information. Our trainable vision projection layer is designed to efficiently aggregate and condense this information.
Let the encoder output a list of $M$ feature tensors. We select a subset of evenly spaced $L$ intermediate layers for processing. For each camera view, the features from a single selected layer $l$ have a dimensionality of $N_{l}\times D_{\text{vggt}}$, where $N_{l}$ is the sequence length of visual tokens and $D_{\text{vggt}}$ is the VGGT’s hidden dimension.
The vision projection layer processes features from the $L$ selected layers for each camera view. Features from each layer ($\mathbf{z}_{l}\in\mathbb{R}^{N_{l}\times D_{\text{vggt}}}$) are first passed through a dedicated 1D convolutional network with trainable convolutional and ReLU layers, followed by an adaptive average pooling layer to pool the features into a single vector ($\mathbf{f}_{l}\in\mathbb{R}^{D_{\text{conv}}}$). These layer-specific vectors are then concatenated into a single vector. This concatenated vector is then passed through a final multi-layer perceptron (MLP) to produce the visual embedding $\mathbf{z}_{\text{vis}}\in\mathbb{R}^{D_{\text{repr}}}$.
The overall transformation for a single view’s features
$\mathbf{z}_{\text{vis}}=\text{MLP}_{\theta}\left(\left[\text{Conv1D}_{\theta_{1}}(\text{Pool}(\mathbf{z}_{l_{1}}));\dots;\right.\right.\\ \left.\left.\text{Conv1D}_{\theta_{L}}(\text{Pool}(\mathbf{z}_{l_{L}}))\right]\right)$
This per-view projection efficiently adapts the powerful, frozen VGGT features using a lightweight, trainable module.
IV-A2 Language and Proprioceptive Encoders
We process non-visual modalities with simple MLP-based projectors.
  • Language Encoder: A task instruction string $l_{t}$ is first encoded using a pretrained sentence transformer [reimers-2019-sentence-bert], yielding an embedding of dimension $D_{\text{lang\_emb}}$. An MLP then projects this embedding into the common representation space, producing $\mathbf{z}_{\text{lang}}\in\mathbb{R}^{D_{\text{repr}}}$.
  • Proprioceptive Encoder: The robot’s state vector $s_{t}$ (end-effector pose and gripper state) is passed through a two-layer MLP to produce the proprioceptive embedding $\mathbf{z}_{\text{proprio}}\in\mathbb{R}^{D_{\text{repr}}}$.
Figure 4: Real-Robot Setup. (a) Illustration of Hardware. (b) The viewpoint used to train all policies. (c) The viewpoint used for our zero-shot evaluation experiments.
IV-B Observation Trunk and Policy Decoder
Our Policy decoder is a GPT-style, decoder-only transformer. The encoded representations from all modalities ($z_{\text{vis}}$ from each camera view, $z_{\text{lang}}$, and $z_{\text{proprio}}$) are treated as a sequence of input tokens. A learnable action token is appended to this sequence. The transformer processes this sequence with a causal self-attention mask, and the output embedding corresponding to the action token’s position, $h_{\text{action}}\in\mathbb{R}^{D_{\text{hidden}}}$, serves as the input to the action head.
IV-C Action Heads
The final action is generated by a dedicated action head module conditioned on the feature vector $h_{\text{action}}$. We experiment with two variantsto handle different action distributions.
  • MLP Head: A simple deterministic head, implemented as a 2-layer MLP, that directly regresses the continuous action vector $a_{t}\in\mathbb{R}^{D_{\text{act}}}$. This head is suitable for unimodal action distributions. we call this variant GeoAware BAKU.
  • VQ-BeT Head: To potentially model multi-modal expert behavior, we use the Vector-Quantized Behavior Transformer (VQ-BeT) head. This module replaces traditional k-means clustering with a VQ-VAE to learn a discrete action codebook. It predicts an action by classifying the appropriate action token from its codebook and regressing a continuous offset, allowing it to capture complex, multi-modal action distributions. This variant is GeoAware VQ-BeT.
Table I: Success Rates (%) on Libero Subsets. Models marked with * are results taken from [kim2025fine].
Libero Subsets Average
Model Spatial Object Goal Long Across Subsets
Original Novel Original Novel Original Novel Original Novel Original Novel
Diffusion Policy* [chi2023diffusion] 78.3 - 92.5 - 68.3 - 50.5 - 72.4 -
Octo* [ghosh2024octo] 78.9 - 85.7 - 84.6 - 51.1 - 75.1 -
OpenVLA-OFT [kim2025fine] 98.0 90.0 98.0 15.7 98.0 81.7 91.0 13.7 96.3 50.2
BAKU [haldar2024baku] 94.0 18.0 100.0 49.3 96.0 80.7 87.0 3.7 94.2 37.9
VQ-BeT [lee2024behavior] 94.0 41.3 100.0 38.3 98.0 83.0 93.0 3.0 96.3 41.4
GeoAware BAKU (ours) 95.0 94.3 98.0 98.0 98.0 90.7 89.9 47.3 95.2 82.6
GeoAware VQ-BeT (ours) 95.0 54.3 100.0 99.0 99.0 85.7 93.0 72.7 96.8 77.9
Table II: Success Rates (%) on Novel Views Across Libero Subsets
Model Small Medium Large Average
OpenVLA-OFT [kim2025fine] 59.3 47 44.5 50.2
BAKU [haldar2024baku] 41.5 31.5 40.8 37.9
VQ-BeT [lee2024behavior] 44.0 43.8 36.5 41.4
GeoAware BAKU 86.0 80.3 81.5 82.6
GeoAware VQ-BeT 88.0 74.5 71.2 77.9
V Experimental Setup
V-A Simulation Benchmarks
The LIBERO benchmark [libero] assesses knowledge transfer in multitask and lifelong robot learning for manipulation tasks. It is composed of four task suites: LIBERO-Spatial, which tests generalization to novel spatial layouts while keeping task and object types fixed; LIBERO-Goal, which introduces new tasks while maintaining the same objects and layouts; LIBERO-Object, which evaluates performance with Novel object types under the same tasks and layouts; and LIBERO-Long, which presents a broader variety of objects, layouts, and backgrounds. Each suite consists of 10 tasks, with 50 human demonstrations provided per task for fine-tuning. We evaluate models trained on Libero’s suites on original and novel viewpoints inspired by [wilcox2025adapt3r] both of which are shown in Figure 3.
V-B Real-World Experiments
To evaluate our approach in the real world, we established a tabletop manipulation environment featuring a Realman 65B robotic arm. We collected a custom dataset by tele-operating the arm using a Meta Quest VR controller. The scene is captured from two fixed, distinct camera viewpoints to provide multi-view observations for the policy. The setup is shown in figure 4.
  1. 1.
    “put both cups on the plate”: Pick up two separate cups and place both of them onto the plate. This is a multi-object manipulation task inspired by Libero Long tasks.
  2. 2.
    “pick the blue cup and place it in the bowl which is on top of the plate”: Pick up the blue cup and place it inside a bowl, which is itself positioned on a plate. This task tests the ability to place an object into a nested target.
  3. 3.
    “pick the pineapple out of the pot and place in between the cups”: This task tests the understanding of spatial relationships by requiring the robot to pick up a pineapple and place it between two other objects.
  4. 4.
    “move the pineapple to the pot and the roll to the plate”: Perform two distinct pick-and-place actions: moving the pineapple to the pot and moving the roll to the plate. This is inspired by tasks from Libero Long.
  5. 5.
    “stack the blue cup on top of the white one”: A stacking cups task which requires high degree of precision for stable vertical alignment especially hard in novel views.
V-C Baselines
We compare our GeoAware-VLA model against three strong baselines: two derived from [haldar2024baku] architecture and OpenVLA-OFT [kim2025fine], which we include to test performance with a larger, pretrained model.
  • BAKU: The standard BAKU model with its original trainable vision encoder and a deterministic MLP action head.
  • VQ-BeT: The BAKU model, but with its MLP head replaced by the VQ-BeT action decoder.
  • OpenVLA-OFT: We utilize the OpenVLA-OFT model as an additional baseline to evaluate the performance of our approach against a larger, more powerful pretrained vision-language model.
The core of our experimental comparison is to isolate the contribution of the geometrically-aware vision encoder. Therefore, the only architectural difference between our GeoAware-VLA models and the corresponding baselines is the vision module. The baselines use the standard trainable vision encoder from BAKU based on Resnet + Film, while our models use the proposed frozen VGGT encoder with a lightweight, trainable projection layer.
Figure 5: (Top) A plot comparing the real-world performance of our GeoAware-VLA against the baseline, highlighting our model’s improvement on both seen and novel viewpoints. (Bottom)The initial and final states for the real-world tasks evaluated.
V-D Training
All models were implemented in PyTorch and trained on a single NVIDIA A100 GPU with 40GB of VRAM. For experiments on the LIBERO benchmark, we trained a separate policy for each of the four task suites. Each model was trained for 150,000 steps using the AdamW optimizer. For our real-world evaluation, policies were trained for 50,000 steps on our custom dataset. Across all experiments, we used a batch size of 64 and provided the policy with observations from two camera views. GeoAware variants were trained with input images resized to $126\times 126$ pixels, whereas the BAKU variants used $128\times 128$ pixels.
VI Results and Discussions
We evaluate GeoAware-VLA to answer four key questions:
  1. 1.
    Is leveraging a frozen, pretrained geometric encoder a viable strategy for imitation learning, even on trained-on views?
  2. 2.
    How effectively does our approach generalize to novel camera viewpoints, and how robust is this generalization to the magnitude of the viewpoint shift?
  3. 3.
    Do the performance gains observed in simulation translate to physical hardware?
  4. 4.
    Which layers from the geometric backbone are most critical for downstream policy performance?
All results are reported as the success rate out of 10 rollouts per task, for both simulation and real-world setups.
VI-A A Geometric Backbone is a Strong In-Distribution Performer
First, we assess whether integrating a frozen geometric backbone affects performance on familiar tasks. As shown in Table I, our GeoAware models not only match the high success rates of their baseline counterparts on the original training views, but also yield an overall improvement in performance, boosting VQ-BeT’s original performance to achieve the highest overall success rate of 96.8%.
This result supports our hypothesis that a lightweight, trainable projection layer is sufficient to adapt the geometrically-rich features from a frozen backbone for the policy.
VI-B GeoAware-VLA Achieves Robust Zero-Shot Generalization Across Viewpoint Shifts
We next evaluate the models’ ability to generalize to novel camera viewpoints in a zero-shot setting. A clear difference in performance emerges when models are evaluated on novel camera poses. The baseline models exhibit a notable decrease in success rates, indicating a sensitivity to viewpoint shifts (Table I).
In contrast, our GeoAware models demonstrate greater robustness, maintaining high success rates and outperforming their baselines by more than 2x, with an overall performance 30% higher than the next baseline. This improvement in generalization highlights the effectiveness of the geometric prior in achieving viewpoint invariance.
To further analyze this robustness, we examine performance across increasing degrees of viewpoint deviation. Table II shows that while baseline performance is lower and inconsistent, our GeoAware models exhibit consistently high success rates and a more robust degradation across all views.
VI-C Real-World Performance Gains
Finally, we evaluate whether the performance improvements observed in simulation translate to a physical system. The results indicate that the trends transfer effectively to the real world. As detailed in Figure 5, our GeoAware BAKU model achieved a measurable performance increase over the baseline on a series of manipulation tasks, even on the original trained viewpoint.
VI-D Ablation on VGGT Layer Selection
To determine which layers from VGGT’s 24-layer backbone are most important for policy performance We compare our default configuration (4 evenly spaced layers) with two alternatives: using all 24 layers, and using only the final 4 layers.
As detailed in Table III, the default configuration using evenly spaced layers achieves the best overall performance. While using all layers provides a small improvement in novel view generalization, it results in a slight drop in in-distribution performance and is more computationally expensive. It is worth noting that even the weakest configuration using the last 4 layer significantly outperforms the next-best baseline on novel views. Despite its overall performance drop, its 34.3% success rate on unseen views is still more than double that of [kim2025fine].
Table III: Ablation on VGGT layer selection for GeoAware VQ-BeT on the LIBERO-Long subset. Success rates (%) are shown for seen and novel viewpoints.
Model Seen Novel
OpenVLA-OFT [kim2025fine] 93.0 13.7
GeoAware VQ-BeT (All Layers) 90.0 74.3
GeoAware VQ-BeT (4 Evenly Spaced - Default) 93.0 72.7
GeoAware VQ-BeT (Last 4 Layers) 60.0 34.3
VII Conclusions and future work
In this paper, we introduced GeoAware-VLA, a method that significantly improves the generalization of VLA models to novel camera viewpoints. By replacing the standard trainable vision encoder with a frozen, pretrained geometric foundation model (VGGT) and a lightweight projection layer, we effectively inject a strong 3D-aware prior into the policy. This simple modification relieves the policy from the difficult task of learning geometric consistency from scratch.
Our experiments on the LIBERO benchmark demonstrate that this approach is highly effective. GeoAware-VLA not only maintains strong performance on the original training viewpoints but also achieves a remarkable improvement of over 2x in zero-shot success rates on novel, unseen views. These gains were consistent across both continuous and discrete action heads and were successfully validated on a real-world robotic manipulation platform.
The success of GeoAware-VLA underscores a key insight: the geometric acuity of the vision backbone is a critical factor for building robust, generalizable robotic agents. Our work presents a practical and computationally efficient blueprint for leveraging powerful, off-the-shelf geometric models to enhance imitation learning policies. Future work could explore integrating other geometric foundation models, applying this method to a broader range of tasks and robot morphologies, and investigating how fine-tuning the VGGT backbone could impact the performance.